OpenStudy (anonymous):

A particle moves along the x-axis with a nonconstant acceleration described by a = 12t, where a is in meters per second squared and t is in seconds. If the particle starts from rest so that its speed v and position x are zero when t = 0, where is it located when t = 2 seconds?

OpenStudy (astrophysics):

You can use calculus right @mathway

OpenStudy (astrophysics):

This will require integration note that \[\vec a = \frac{ d^2x }{ dt^2 } = 12t\]

OpenStudy (astrophysics):

Or we can use \[a = \frac{ dv }{ dt }\] so \[\int\limits dv = \int\limits 12t dt\]

OpenStudy (anonymous):

I hope so, but this is AP Physics 1, and it is Algebra-based.